I have made 60 cupcakes. 36 are chocolate and 24 are vanilla. How many more chocolate do I need to make to have 70% of the cupcakes be chocolate. Please show the proportion used.

Answer by stanbon(75887)

I have made 60 cupcakes. 36 are chocolate and 24 are vanilla. How many more chocolate do I need to make to have 70% of the cupcakes be chocolate. Please show the proportion used.
-----
Proportion:
(# of chocolate)/(total # of cupcakes) = 70/100
-------
(36+x)/(60+x) = 70/100
----
36*100 + 100x = 70*60 + 70x
-----
30x = 600
x = 20 (# of chocolate to add)
============
Cheers,
Stan H.